# 6G NTN Waveforms: A Comparison of OTFS, AFDM and OCDM in LEO Satellite Channels

### v1

@@ I Introduction
Non-terrestrial networks (NTNs) have been an integral part of the 3rd generation partnership project (3GPP) releases from release 17 (R17) onwards for standardizing 5th generation (5G) new radio (NR) [18]. In 6G, whose study phase is marked by R20, NTNs are expected to play a key role in achieving ubiquity, transparency, reconfigurability, and resilience through NTN-Internet of Things (IoT) coalescence. There are a plethora of NTN nodes ranging from different types of satellites to high-altitude platforms (HAPs) and to unmanned aerial vehicles (UAVs). However, thanks to the recent agreements between commercial satellite operators and mobile network operators (e.g. Starlink-TMobile direct to cell service), low Earth orbit (LEO) satellites have become native to the cellular fabric. No wonder R20 is provisioning special study items for LEO satellites, including seamless handover between gNodeB and LEO satellite, Doppler/ high-velocity considerations, and synchronization without global navigation satellite systems (GNSS) [4]. A typical NTN architecture with a LEO satellite and 3GPP interfaces [6] for highly mobile scenarios is shown in Fig. 1.
@@ II Waveform Model
$\mathbf{s}_{\mathrm{ofdm}}=\mathbf{F}_{N}^{-1}\cdot\mathbf{x},$ (1)
where $\mathbf{F}_{N}\in\mathbb{C}^{N\times N}$ is a $N$-point FFT matrix, and its inverse, $\mathbf{F}_{N}^{-1}\triangleq\mathbf{F}_{N}^{\text{H}}$, is a $N$-point IFFT matrix [12]. The noise corrupted version at receiver is given by,
$\mathbf{r}_{\mathrm{ofdm}}=\mathbf{H}\cdot\mathbf{s}_{\mathrm{ofdm}}+\mathbf{w},$ (2)
where $\mathbf{H}$ and $\mathbf{w}$ denote channel matrix and noise matrix, respectively. The symbols recoverd at receiver end, $\mathbf{y}_{\text{ofdm}}\in\mathbb{C}^{N\times 1}$, making use of $N$-point FFT is:
$\mathbf{y}_{\mathrm{ofdm}}=\mathbf{F}_{N}\cdot\mathbf{r}_{\text{ofdm}}={(\mathbf{F}_{N}\cdot\mathbf{H}\cdot\mathbf{F}_{N}^{\text{H}})}\cdot\mathbf{x}+\mathbf{F}_{N}\cdot\mathbf{w}.$ (3)
@@
$\mathbf{H}_{\mathrm{eff}}^{\mathrm{ofdm}}=\mathbf{F}_{N}\cdot\mathbf{H}\cdot\mathbf{F}_{N}^{\text{H}}.$ (4)
@@ II-A AFDM
$\mathbf{s}_{\mathrm{afdm}}=(\mathbf{\Lambda}_{c_{1}}\cdot\mathbf{F}_{N}\cdot\mathbf{\Lambda}_{c_{2}})^{-1}\cdot\mathbf{x}=\mathbf{\Lambda}_{c_{2}}^{\mathrm{H}}\cdot\mathbf{F}_{N}^{\mathrm{H}}\cdot\mathbf{\Lambda}_{c_{1}}^{\mathrm{H}}\cdot\mathbf{x},$ (5)
@@
$\mathbf{y}_{\mathrm{afdm}}=\mathbf{DAFT}\cdot(\mathbf{H}\cdot\mathbf{s}_{\text{afdm}}+\mathbf{w})=\mathbf{H}_{\mathrm{eff}}^{\text{afdm}}\cdot\mathbf{x}+\mathbf{DAFT}\cdot\mathbf{w},$ (7)
where, $\mathbf{DAFT}\triangleq\mathbf{\Lambda}_{c_{1}}\cdot\mathbf{F}_{N}\cdot\mathbf{\Lambda}_{c_{2}}$, whereas $\mathbf{H}$ and $\mathbf{H}_{\mathrm{eff}}^{\text{afdm}}$ are the doubly-dispersive channel matrix and effective channel matrix for AFDM, respectively.
@@ II-C OTFS
$\mathbf{s}_{\mathrm{otfs}}$ $=\operatorname{vec}\left(\mathbf{P}^{\mathrm{tx}}\cdot\mathbf{F}_{K}^{-1}\left(\mathbf{F}_{K}\cdot\mathbf{x}\cdot\mathbf{F}_{L}^{-1}\right)\right)$
@@
$\mathbf{y}_{\mathrm{otfs}}$ $=\left(\mathbf{F}_{L}\otimes\mathbf{P}^{\mathrm{rx}}\right)\left(\mathbf{H}\cdot\mathbf{s}_{\mathrm{otfs}}+\mathbf{w}\right)$
$=\mathbf{H}_{\mathrm{eff}}^{\mathrm{otfs}}\cdot\mathbf{x}+\left(\mathbf{F}_{L}\otimes\mathbf{P}^{\mathrm{rx}}\right)\mathbf{w},$ (9)

### v2

@@ I Introduction
Non-terrestrial networks (NTNs) have been an integral part of the 3rd generation partnership project (3GPP) releases from release 17 (R17) onwards for standardizing 5th generation (5G) new radio (NR) [18]. In 6G, whose study phase is marked by R20, NTNs are expected to play a key role in achieving ubiquity, transparency, reconfigurability, and resilience through NTN-Internet of Things (IoT) coalescence. There are a plethora of NTN nodes ranging from different types of satellites to high-altitude platforms (HAPs) and to unmanned aerial vehicles (UAVs). However, thanks to the recent agreements between commercial satellite operators and mobile network operators (e.g., Starlink T-Mobile direct handset to satellite service), low Earth orbit (LEO) satellites have become native to the cellular fabric. No wonder R20 is provisioning special study items for LEO satellites, including seamless handover between gNodeB and LEO satellite, Doppler/ high-velocity considerations, and synchronization without global navigation satellite systems (GNSS) [4]. A typical NTN architecture with a LEO satellite and 3GPP interfaces [6] for highly mobile scenarios is shown in Fig. 1.
@@ II Waveform Model
$\mathbf{s}_{\mathrm{ofdm}}=\mathbf{F}_{N}^{\mathrm{H}}\mathbf{x},$ (1)
where $\mathbf{F}_{N}\in\mathbb{C}^{N\times N}$ is a $N$-point FFT matrix, and its conjugate transpose, $\mathbf{F}_{N}^{\text{H}}$, is a $N$-point IFFT matrix [12]. The noise corrupted version at receiver is given by,
$\mathbf{r}_{\mathrm{ofdm}}=\mathbf{H}\mathbf{s}_{\mathrm{ofdm}}+\mathbf{w},$ (2)
where $\mathbf{H}$ and $\mathbf{w}$ denote channel matrix and noise matrix, respectively. The symbols recoverd at receiver end, $\mathbf{y}_{\mathrm{ofdm}}\in\mathbb{C}^{N\times 1}$, making use of $N$-point FFT is:
$\mathbf{y}_{\mathrm{ofdm}}=\mathbf{F}_{N}\mathbf{r}_{\text{ofdm}}={(\mathbf{F}_{N}\mathbf{H}\mathbf{F}_{N}^{\mathrm{H}})}\mathbf{x}+\mathbf{F}_{N}\mathbf{w}.$ (3)
@@
$\mathbf{H}_{\mathrm{eff}}^{\mathrm{ofdm}}=\mathbf{F}_{N}\mathbf{H}\mathbf{F}_{N}^{\mathrm{H}}.$ (4)
@@ II-A AFDM
$\mathbf{s}_{\mathrm{afdm}}=(\mathbf{\Lambda}_{c_{1}}\mathbf{F}_{N}\mathbf{\Lambda}_{c_{2}})^{\mathrm{H}}\mathbf{x}=\mathbf{\Lambda}_{c_{2}}^{\mathrm{H}}\mathbf{F}_{N}^{\mathrm{H}}\mathbf{\Lambda}_{c_{1}}^{\mathrm{H}}\mathbf{x},$ (5)
@@
$\mathbf{y}_{\mathrm{afdm}}=\mathbf{DAFT}\cdot(\mathbf{H}\mathbf{s}_{\text{afdm}}+\mathbf{w})=\mathbf{H}_{\mathrm{eff}}^{\text{afdm}}\mathbf{x}+\mathbf{DAFT}\cdot\mathbf{w},$ (7)
where, $\mathbf{DAFT}\triangleq\mathbf{\Lambda}_{c_{1}}\mathbf{F}_{N}\mathbf{\Lambda}_{c_{2}}$, whereas $\mathbf{H}$ and $\mathbf{H}_{\mathrm{eff}}^{\text{afdm}}$ are the doubly-dispersive channel matrix and effective channel matrix for AFDM, respectively.
@@ II-C OTFS
$\mathbf{s}_{\mathrm{otfs}}$ $=\operatorname{vec}\left(\mathbf{P}^{\mathrm{tx}}\mathbf{F}_{K}^{-1}\left(\mathbf{F}_{K}\mathbf{x}\mathbf{F}_{L}^{-1}\right)\right)$
@@
$\mathbf{y}_{\mathrm{otfs}}$ $=\left(\mathbf{F}_{L}\otimes\mathbf{P}^{\mathrm{rx}}\right)\left(\mathbf{H}\mathbf{s}_{\mathrm{otfs}}+\mathbf{w}\right)$
$=\mathbf{H}_{\mathrm{eff}}^{\mathrm{otfs}}\mathbf{x}+\left(\mathbf{F}_{L}\otimes\mathbf{P}^{\mathrm{rx}}\right)\mathbf{w},$ (9)